What is the partial fraction decomposition of x/(x+3)(x-1)
Answer by stanbon(75887)

What is the partial fraction decomposition of x/(x+3)(x-1)
-------
x/[(x+3)(x-1)] = A/(x+3) + B/(x-1)
----------------------------------------
= [A(x-1)+B(x+3)]/(x+3)(x-1)
-------
Equation the numerators:
x = (A+B)x + (3B-A)
------
Equations:
A + B = 1
-A+3B = 0
-----------
Add and solve for "B":
4B = 1
B = 1/4
--------
Solve for "A":
A + (1/4) = 1
A = 3/4
--------
Partial-fraction decomposition: (3/4)/(x+3) + (1/4)/(x-1)
=========================================================
Cheers,
Stan H.